Question 1 1 pts Carneys current income statement shows the following: $500,000 Sales revenue Cost of goods sold Variable coQuestion 2 1 pts Truman Company has developed a new gadget with the following cost information Variable manufacturing costs:Question 3 1 pts The following data is available for HVAC Specialists: Repair specialists wages $ 150,000 Repair specialists
Answer #1

1.

Contribution margin = Sales price - Variable costs

= 500,000 - (150,000+100,000)

= 250,000

Contribution margin ratio = 250,000/500,000 = 50%

Breakeven sales = Fixed cost/CM ratio

= (150,000+50,000)/50% = 400,000

Margin of Safety = Sales - Breakeven sales

= 500,000 - 400,000

= 100,000

Option B